# Mitigating the Impact of Outlier Channels for Language Model Quantization with Activation Regularization

### v1

@@ Abstract
3IBM Research 4IST Austria 5NeuralMagic
@@ 1 Introduction
Large language models (LLM) have been shown to contain outlier channels, i.e., feature dimensions whose values are orders of magnitude higher than the others. These outlier channels are known to be crucial for strong model performance (Kovaleva et al., 2021; Puccetti et al., 2022), but pose significant challenges from a model compression perspective, for instance via post-training quantization (PTQ) (Dettmers et al., 2022; Xiao et al., 2023; Wei et al., 2022). Concretely, to enable the use of low-bitwidth integer matrix multiplications—which can lead to significant speed-ups—both the activations and the weights need to be quantized. However the presence of high outlier values in the model activations results in high quantization errors, and thus overall poor PTQ accuracy (see, e.g., Xiao et al. (2023)).
@@ 3 Motivating Study: Outlier Channels in Language Models
Figure 5: Activation development in two open-source models: Pythia 6.9b (Biderman et al., 2023) and OLMo 7B (Groeneveld et al., 2024). We show activiations for a layer that reads from the residual stream (QKV Input) and one that does not (MLP Proj Input). Note that the OLMo data includes a step-0 checkpoint (i.e., at initialization).
@@ 4.2 Output Activations: Kurtosis Regularization
One approach to mitigating these outlier weights would be to directly regularize the weights via QAT or some other approach (e.g., $\ell_{\infty}$-norm regularization). However, we found these direct regularization approaches to result in much worse performance and/or unstable training. We thus adopt a more indirect regularization strategy, exploiting the fact that high input channel weights typically lead to a layer’s outputs having outliers, i.e., the output distribution is heavy-tailed (see fig. 1). Our approach thus regularizes the output distribution’s kurtosis. which measures how heavy-tailed a distribution is. An estimate of the kurtosis of a set of values $\mathbf{x}\in\mathbb{R}^{d}$ is given by, $\operatorname{Kurtosis}(\mathbf{d})=\frac{\sum_{i}^{k}(\mathbf{x}_{i}-\mu)^{4}}{\sigma^{4}+\epsilon},$ where $\mu$ and $\sigma$ are respectively the empirical mean and standard deviation of $\mathbf{x}$, and $\epsilon$ is a small term for numerical stability. We multiply the sum of the kurtosis estimates for each token with hyperparameter $\lambda$, and add the result to the cross-entropy loss. While prior work has shown the benefits of regularizing the kurtosis of a layer’s activation distribution to be close to that of a uniform distribution (Chmiel et al., 2020), regularizing the output distribution’s kurtosis to make it less heavy-tailed has not been explored before to our knowledge.
@@ 5.2 Results
We report the results of our 1B experiments on the C4 and PTB dataset in table 1. We observe that our approach can learn a W4A4 model that has respectable performance compared to the W16A16 baseline. We also observe that the gap between the QAT model with and without kurtosis expands as weights are quantized more and more. At full precision, the gap is less than 1%. At 4 bits, this expands to between 3% and 4%, and at 3 bits this gap widens to 21%. All non-QAT method have catastrophic performance degradations with 4-bit activations. Activation clamping is the only method that achieves less than two orders of magnitude increase in perplexity. In table 2 we perform experiments on downstream tasks for select models to validate our usage of perplexity as a proxy for downstream performance. We observe that models with similar perplexity exhibit similar downstream performance.
@@ Acknowledgments
This study was supported by funds from an MIT-IBM Watson AI grant.

### v2

@@ Abstract
3IBM Research 4IST Austria 5Neural Magic, Inc.
@@ 1 Introduction
Large language models (LLM) have been shown to contain outlier channels, i.e., feature dimensions whose values are orders of magnitude higher than the others. These outlier channels are known to be highly correlated with strong model performance (Kovaleva et al., 2021; Puccetti et al., 2022), but pose significant challenges from a model compression perspective, for instance via post-training quantization (PTQ) (Dettmers et al., 2022; Xiao et al., 2023; Wei et al., 2022). Concretely, to enable the use of low-bitwidth integer matrix multiplications—which can lead to significant speed-ups—both the activations and the weights need to be quantized. However the presence of high outlier values in the model activations results in high quantization errors, and thus overall poor PTQ accuracy (see, e.g., Xiao et al. (2023)).
@@ 3 Motivating Study: Outlier Channels in Language Models
Figure 5: Activation development in two open-source models: Pythia 6.9b (Biderman et al., 2023) and OLMo 7B (Groeneveld et al., 2024). We show activations for a layer that reads from the residual stream (QKV Input) and one that does not (MLP Proj Input). Note that the OLMo data includes a step-0 checkpoint (i.e., at initialization).
@@ 4.2 Output Activations: Kurtosis Regularization
One approach to mitigating these outlier weights would be to directly regularize the weights via QAT or some other approach (e.g., $\ell_{\infty}$-norm regularization). However, we found these direct regularization approaches to result in much worse performance and/or unstable training. We thus adopt a more indirect regularization strategy, exploiting the fact that high input channel weights typically lead to a layer’s outputs having outliers, i.e., the output distribution is heavy-tailed (see fig. 1). Our approach thus regularizes the output distribution’s kurtosis, which intuitively measures how heavy-tailed a distribution is. An estimate of the kurtosis of a set of values $\mathbf{x}\in\mathbb{R}^{d}$ is given by
$\operatorname{Kurtosis}(\mathbf{d})=\frac{\sum_{i}^{k}(\mathbf{x}_{i}-\mu)^{4}}{\sigma^{4}+\epsilon},$
where $\mu$ and $\sigma$ are respectively the empirical mean and standard deviation of $\mathbf{x}$, and $\epsilon$ is a small term for numerical stability. We multiply the sum of the kurtosis estimates for each token with hyperparameter $\lambda$, and add the result to the cross-entropy loss. While prior work has shown the benefits of regularizing the kurtosis of a layer’s activation distribution to be close to that of a uniform distribution (Chmiel et al., 2020), regularizing the output distribution’s kurtosis to make it less heavy-tailed has not been explored before to our knowledge.
@@ 5.2 Results
We report the results of our 1B experiments on the C4 and PTB dataset in table 1. We observe that our approach can learn a W4A4 model that has respectable performance compared to the W16A16 baseline. We also observe that the gap between the QAT model with and without kurtosis expands as weights are quantized more and more. At full precision, the gap is less than 1%. At 4 bits, this expands to between 3% and 4%, and at 3 bits this gap widens to 21%. All non-QAT methods have catastrophic performance degradations with 4-bit activations: Activation clamping is the only method that achieves less than two orders of magnitude increase in perplexity. In table 2 we perform experiments on downstream tasks for select models to validate our usage of perplexity as a proxy for downstream performance. We observe that models with similar perplexity exhibit similar downstream performance.
@@ Acknowledgments
This study was supported by MIT-IBM Watson AI Lab.